OpenStudy (anonymous):

6/5 multiplied by 12

OpenStudy (anonymous):

14.4

OpenStudy (anonymous):

\[(\frac{ 6 }{ 5 })(\frac{ 60 }{ 5 }) = \frac{ 360 }{ 25 } = \frac{ 72 }{ 5 }\] 60/5 is 12

OpenStudy (anonymous):

6/5(12) ??

OpenStudy (anonymous):

14.4

OpenStudy (nincompoop):

that's a long way to go and I do not know why you have to go through the length of converting 12 into 60/5 \[ \sf \huge \frac{6}{5} \times \frac{12}{1} \]

OpenStudy (dmndlife24):

6/5(12) = 72/5 or 14.4 :)

OpenStudy (anonymous):

cant be a decimal

OpenStudy (dmndlife24):

Well...it is

OpenStudy (dmndlife24):

why CAN't it be??

OpenStudy (anonymous):

WebWork. Says I cant use a decimal

OpenStudy (nincompoop):

do you know how to multiply and divide?

OpenStudy (anonymous):

y-14=-6/5(x+12) put in slope intercept form

OpenStudy (nincompoop):

what the f#%ck

OpenStudy (nincompoop):

that is completely different from what you posted

OpenStudy (dmndlife24):

Woah...WUT?

OpenStudy (anonymous):

don't you have to multiply 6/5 by 12 anyways?

OpenStudy (anonymous):

What you have their now is an equation of a line. What do you have to do with it?

OpenStudy (dmndlife24):

y=-6/5x + 142/5

OpenStudy (anonymous):

says its wrong

OpenStudy (dmndlife24):

WHAT is wrong??

OpenStudy (anonymous):

-6/5x+142/5

OpenStudy (anonymous):

Still do not understand. What is the question specifically that you have to do with this line equation?

OpenStudy (anonymous):

just solve for y?

OpenStudy (dmndlife24):

It's in slope-int. form just like you needed...

OpenStudy (anonymous):

I dont understand either

OpenStudy (anonymous):

well what is the entire question or can you take a screenshot of your screen for us?

OpenStudy (anonymous):

how do i post it to here

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

does anyone know? :(

OpenStudy (anonymous):

You y intercept is wrong in slope-intercept form \[y-14 = \frac{ -6 }{ 5 }(x + 12)\] \[y-14 = \frac{ -6 }{ 5 }x - \frac{ 36 }{ 5 }\] Get 14 to have a common denominator of 5 which is 70/5 \[y - \frac{ 70 }{ 5 } = \frac{ -6 }{ 5 }x - \frac{ 36 }{ 5 }\] \[y = \frac{ -6 }{ 5 }x + \frac{ 34 }{ 5 }\]

OpenStudy (anonymous):

ahhh woops lol i multipled -6 and 12 wrong

OpenStudy (anonymous):

Hang on let me redo

OpenStudy (anonymous):

\[y - 14 = \frac{ -6 }{ 5 }x - \frac{ 72 }{ 5 }\] \[y - \frac{ 70 }{ 5 } = \frac{ -6 }{ 5 }x - \frac{ 72 }{ 5 }\] \[y = \frac{ -6 }{ 5 }x - \frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem